OpenStudy (cherokeegutierrez):

WILL GIVE MEDAL! What is the value of 2x^-2 y^-2 for x=3 and y=-2

OpenStudy (rushwr):

\[2x ^{-2}y ^{-2}\] this what u have said right? Just substitute the values and see

OpenStudy (cherokeegutierrez):

yes. but i need help

OpenStudy (rushwr):

which part do u need help

OpenStudy (cherokeegutierrez):

figuring the whole thing out

OpenStudy (rushwr):

okai, \[2x ^{-2}y ^{-2} = 2 * 3^{-2} * -2^{-2}\] right?

OpenStudy (cherokeegutierrez):

i guess idk thats why i need help

OpenStudy (rushwr):

in tha step I have substituted the values of x and y which we are given with ! Did u get that part ?

OpenStudy (cherokeegutierrez):

yes.

OpenStudy (rushwr):

When u have a negative exponent like \[x ^{-n}\] You can also write it as \[\frac{ 1 }{ x ^{n} }\] okai ?

OpenStudy (cherokeegutierrez):

ok

OpenStudy (cherokeegutierrez):

ok so what now?

OpenStudy (rushwr):

I got a little wrong there y also has a negative exponent so we can take that down too ! so it will be \[\frac{ 2 }{ 3^{2} * -2^{2} }\]

OpenStudy (cherokeegutierrez):

ok lol i was like why did you dlete it. but anyways i got it. whats next

OpenStudy (rushwr):

now just simplyfy when u multiply 2 negatives it becomes positive right? we knwo -2 ^2 means -2 x -2 right? so that makes it 4!

OpenStudy (cherokeegutierrez):

so 2/2?

OpenStudy (rushwr):

we know 3^2 = 9 \[\frac{ 2 }{ 9 * 4 }\]

OpenStudy (rushwr):

did u get it ?

OpenStudy (cherokeegutierrez):

so 2/36

OpenStudy (cherokeegutierrez):

but there is no answer like that on here

OpenStudy (cherokeegutierrez):

ohhh 1/18?

OpenStudy (cherokeegutierrez):

there is that one